Safrole is contained in oil of sassafras and was once used to flavor root beer. A 2.39-mg sample of safrole was dissolved in 103.0 mg of diphenyl ether. The solution had a melting point of 25.70°C. Calculate the molecular mass of safrole. The freezing point of pure diphenyl ether is 26.84°C, and the freezing-point depression constant, Kf, is 8.00°C/m.


1
Expert's answer
2020-11-02T09:02:44-0500

∆T=KfCm, where Cm is the molality of the solution.

∆T=T0-Tf=26.84-25.7=1.14°C,

Cm=∆T/Kf=1.14/8.00=0.1425(mol/kg),

Cm=nsolute/msolvent

nsolute=Cm•msolvent=0.1425×0.000103=0.0000147(mol)

M=m/n=0.00239/0.0000147=162.59(g/mol)